OpenStudy (anonymous):

Bernoulli differential equation by using substitution 3(1+t^2) dy/dx = 2ty(y^3 - 1) Plz help. Will greatly appreciate the detailed answer.

OpenStudy (anonymous):

\[\begin{align*} 3(1+t^2)\frac{dy}{dt}&=2ty(y^3-1)\\\\ 3(1+t^2)\frac{dy}{dt}&=2ty^4-2ty\\\\ 3(1+t^2)y^{-4}\frac{dy}{dt}&=2t-2ty^{-3} \end{align*}\] Set \(u=y^{-3}\), then \(\dfrac{du}{dt}=-3y^{-4}\). Then \[\begin{align*} -(1+t^2)\frac{du}{dt}&=2t-2tu\\\\ \frac{du}{dt}-\frac{2t}{1+t^2}u&=\frac{2t}{1+t^2} \end{align*}\] The equation is now linear in \(u\).